# uncompress

### v2

@@ 4.4. Proof of bounded irrationality of irreducible fibres
Recall that there is a birational $Z$-map $\phi\colon X\dashrightarrow Y$ and that $F$ is the reduction of the central fibre of $X$ over the closed point $z\in Z$. Moreover, by the construction in Lemma 4.3, $\phi$ does not contract any horizontal/Z/Z divisors, and all the horizontal/Z/Z exceptional divisors of $\phi^{-1}$ are contained in $\operatorname{Supp}D$. Furthermore, by Theorem 3.12 (ii), $m_{Y}^{-1}(\operatorname{Supp}D)$ is contained in $Y^{\prime}\setminus U_{Y^{\prime}}$.
Step 3. By Lemma 3.7 and the choice of the closed point $z\in Z$, the reduced divisor $\{z\}$ and the support of the fibre of $g^{\prime}$ over $z$ are contained in the toroidal boundaries of $(U_{Z^{\prime}}\subset Z^{\prime})$ and $(U_{Y^{\prime}}\subset Y^{\prime})$ respectively. Denote by $D^{\prime}$ the reduced divisor supporting on the complement $Y^{\prime}\setminus U_{Y^{\prime}}$. Define the divisor $D_{Y^{\prime}}$ by $K_{Y^{\prime}}+D_{Y^{\prime}}=(\phi^{-1}\circ m_{Y})^{*}(K_{X}+B^{+})$. Similarly, denote by $D_{Y}$ the divisor defined by $K_{Y}+D_{Y}=(\phi^{-1})^{*}(K_{X}+B^{+})$. Note that $D_{Y^{\prime}}$ and $D_{Y}$ can have components with negative coefficients. Since $Y$ is generically smooth over $Z$, over the generic point of $Z$, we can write
$K_{Y^{\prime}}+R=m_{Y}^{*}K_{Y}$
where $R\leq 0$ supports in the exceptional locus of $m_{Y}$, and hence we have
@@
over the generic point of $Z$. By the construction of the divisor $D$ on $Y$ in Lemma 4.3, $\operatorname{Supp}D_{Y}$ is contained in $\operatorname{Supp}D$ over the generic point of $Z$. Moreover, since $m_{Y}^{-1}(\operatorname{Supp}D)$ is contained in $\operatorname{Supp}D^{\prime}$ by construction, we conclude that $\operatorname{Supp}m_{Y}^{*}D_{Y}$ is contained in $\operatorname{Supp}D^{\prime}$ over the generic point of $Z$. Then as $D_{Y^{\prime}}\leq 1$, we have $D_{Y^{\prime}}\leq D^{\prime}$ over the generic point of $Z$. On the other hand, as the support of the fibre of $g^{\prime}$ over $z\in Z$ is contained in the toroidal boundary $D^{\prime}$, shrinking $Z$ around $z$ if necessary, we can assume that $D_{Y^{\prime}}\leq D^{\prime}$ over the whole $Z$.
@@
Step 5. Now let $V$ be an lc centre of $(Y^{\prime},D^{\prime})$ that is contained in $S$. We show that $V$ is birationally bounded. Notice that a general fibre of $\Phi^{\prime}$ is normal, hence, over the generic point of $Z$, the normal scheme $Y^{\prime}$ is isomorphic to the fibre product $\mathfrak{X}^{\prime}\times_{\mathfrak{B}^{\prime}}Z^{\prime}$. Thus, by the construction in Step 3, there is an integer $r\in\mathbb{N}$ such that the volume of $K_{Y^{\prime}}+D^{\prime}+A$ on a general fibre of $Y^{\prime}\to Z^{\prime}$ is less than $r$. Let $W$ be the normalisation of $V$. By [FH23, Theorem 1.1, Definition 1.3], we can write $(K_{Y^{\prime}}+D^{\prime})|_{W}=K_{W}+D_{W}+M_{W}$ where $(W,D_{W}+M_{W})$ is an NQC generalised lc generalised pair. Now let $L=K_{Y^{\prime}}+D^{\prime}+A$, and let $L_{W}=L|_{W}$ which is an ample Cartier divisor on $W$ as $L$ is ample and Cartier. By applying Lemma 2.11 and its proof to $L_{W}=K_{W}+D_{W}+M_{W}+A_{W}$, we see that $\left|mL_{W}\right|$ is birational for some $m\in\mathbb{N}$ depending only on $\dim W$. It is then enough to show that the volume of $L_{W}$ is bounded by [HMX13, Lemma 2.4.2 (2)].
Pick a component $S^{\prime}$ of $S$ containing $V$ and let $T$ be its normalisation. Applying [FH23, Theorem 1.1], we can write $(K_{Y^{\prime}}+D^{\prime})|_{T}=K_{T}+D_{T}$ where $(T,D_{T})$ is lc (since $T$ is codimension one there is no nef part of the generalised pair in this adjunction). Take a general Cartier divisor $P\geq 0$ on $Y^{\prime}$ containing $V$ and avoiding other lc centres of $(Y^{\prime},D^{\prime})$. For any small $\alpha>0$, $(Y^{\prime},D^{\prime}+\alpha P)$ is not lc near $V$, hence applying [FH23, Theorem 1.1] again we deduce that $(T,D_{T})$ has an lc centre $V_{T}$ mapping onto $V$.
Denote by $A_{T}$ the pullback of the ample Cartier divisor $A$ to $T$. By assumption on volumes, we have that the volume of the ample Cartier divisor $K_{T}+D_{T}+A_{T}=(K_{Y^{\prime}}+D^{\prime}+A)|_{T}$ is $\leq r$, so the volume of $K_{T}+D_{T}+A_{T}$ takes only finitely many values. Changing the ample Cartier divisor $A_{T}$ linearly, we can assume that $(T,D_{T}+A_{T})$ is lc. Then [HMX18, Theorem 1.1] shows that $(T,D_{T}+A_{T})$ belongs to a bounded set of pairs, so the lc centre $V_{T}$ of $(T,D_{T})$ is also bounded by Lemma 2.10. This implies that the volume of $(K_{Y^{\prime}}+D^{\prime}+A)|_{V_{T}}$ is bounded, hence the volume of $(K_{Y^{\prime}}+D^{\prime}+A)|_{W}$ is bounded as desired.
Step 6. Now we can conclude the proof as follows. Recall that $\operatorname{centre}_{Y^{\prime}}(F)$ is an lc centre of the pair $(Y^{\prime},D^{\prime})$ that is contained in $S$. By Step 5, $\operatorname{centre}_{Y^{\prime}}(F)$ is birationally bounded. Denote by $C$ a bounded nonsingular resolution of $\operatorname{centre}_{Y^{\prime}}(F)$, hence the irrationality $\operatorname{irr}(C)$ is bounded from above by Lemma 2.17. Moreover, by Proposition 3.5, $F\to\operatorname{centre}_{Y^{\prime}}(F)$ has irreducible and rational general fibres, so does the rational map $F\dashrightarrow C$. ∎
@@
Let $F$ be an irreducible component of a closed fibre of $f$. Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety hence with bounded degree of irrationality.
@@
Let $z\in Z$ be the closed point so that $F$ is contained in the fibre $f^{-1}(z)$. By [BCHM10], we can run $\phi\colon X\dashrightarrow X^{\prime}$ a (F)(-F)-MMP over $Z$ that ends with a good minimal model of F-F. By negativity lemma, $F$ is not contracted by $\phi$. Let $F^{\prime}\coloneqq\phi_{*}F$. Then $F^{\prime}$ supports on the whole fibre of $X^{\prime}\to Z$ over $z$. Note that $K_{X^{\prime}}$ is ample over the generic point of $Z$, so $K_{X^{\prime}}$ is big/Z/Z. Again by [BCHM10], we can run $\varphi\colon X^{\prime}\dashrightarrow X^{\prime\prime}$ a $(-K_{X^{\prime}})$-MMP over $Z$ that ends with a Fano fibration $g\colon X^{\prime\prime}\to Z$. As $F^{\prime}$ supports on the whole fibre of $X^{\prime}\to Z$ over $z$, we see that $F^{\prime\prime}\coloneqq\varphi_{*}F^{\prime}$ is an irreducible divisor supporting on the whole fibre of $g$ over $z$. Up to decreasing $\epsilon$ slightly, we can take an effective $\mathbb{Q}$-divisor $B$ so that $(X,B)$ is $\epsilon$-lc, and $K_{X}+B\sim_{\mathbb{Q}}0/Z$. Let $B^{\prime\prime}$ be the pushdown of $B$ to $X^{\prime\prime}$. As $\phi,\varphi$ are birational contractions, $(X^{\prime\prime},B^{\prime\prime})$ is also $\epsilon$-lc, and $K_{X^{\prime\prime}}+B^{\prime\prime}\sim_{\mathbb{Q}}0/Z$. By [Bir23a, Theorem 1.1], the lc threshold of $g^{*}z$ with respect to $(X^{\prime\prime},B^{\prime\prime})$ is bounded from below. Then there exists a $t>0$ depending only on $d,\epsilon$ such that $(X^{\prime\prime},B^{\prime\prime}+tF^{\prime\prime})$ is lc, hence $(X^{\prime\prime},tF^{\prime\prime})$ is lc. Thus, we can conclude by Theorem 1.3. ∎

### v3

@@ 4.4. Proof of bounded irrationality of irreducible fibres
Recall that there is a birational map $\phi\colon X\dashrightarrow Y$ over $Z$ and that $F$ is the reduction of the closed fibre of $X\to Z$ over the closed point $z\in Z$. Moreover, by the construction in Lemma 4.3, $\phi$ does not contract any horizontal/Z/Z divisors, and all the horizontal/Z/Z exceptional divisors of $\phi^{-1}$ are contained in $\operatorname{Supp}D$. Furthermore, by Theorem 3.12 (ii), $m_{Y}^{-1}(\operatorname{Supp}D)$ is contained in $Y^{\prime}\setminus U_{Y^{\prime}}$.
Step 3. By Lemma 3.7 and the choice of the closed point $z\in Z$, the reduced divisor $\{z\}$ and the support of the fibre of $g^{\prime}$ over $z$ are contained in the toroidal boundaries of $(U_{Z^{\prime}}\subset Z^{\prime})$ and $(U_{Y^{\prime}}\subset Y^{\prime})$ respectively. Denote by $D^{\prime}$ the reduced divisor supported on the complement $Y^{\prime}\setminus U_{Y^{\prime}}$. Define the $\mathbb{Q}$-divisor $D_{Y^{\prime}}$ by $K_{Y^{\prime}}+D_{Y^{\prime}}=(\phi^{-1}\circ m_{Y})^{*}(K_{X}+B^{+})$. Similarly, denote by $D_{Y}$ the $\mathbb{Q}$-divisor defined by $K_{Y}+D_{Y}=(\phi^{-1})^{*}(K_{X}+B^{+})$. Note that $D_{Y^{\prime}}$ and $D_{Y}$ may have irreducible components with negative coefficients. Since $Y$ is generically smooth over $Z$, over the generic point of $Z$, we can write
$K_{Y^{\prime}}+R=m_{Y}^{*}K_{Y},$
where $R\leq 0$ is supported in the exceptional locus of $m_{Y}$, hence we have
@@
over the generic point of $Z$. By construction of the reduced divisor $D$ on $Y$ in Lemma 4.3, $\operatorname{Supp}D_{Y}$ is contained in $\operatorname{Supp}D$ over the generic point of $Z$. Moreover, since $m_{\mathfrak{X}}^{-1}(\operatorname{Supp}\mathfrak{D})$ is contained in $\mathfrak{X}^{\prime}\setminus U_{\mathfrak{X}^{\prime}}$, and since $U_{Y^{\prime}}=h_{Y^{\prime}}^{-1}(U_{\mathfrak{X}^{\prime}})$ and $D=\mathfrak{D}\times_{\mathfrak{B}}Z$, by the commutative diagram in Step 2, $m_{Y}^{-1}(\operatorname{Supp}D)$ is contained in $\operatorname{Supp}D^{\prime}$. Hence we can conclude that $\operatorname{Supp}m_{Y}^{*}D_{Y}$ is contained in $\operatorname{Supp}D^{\prime}$ over the generic point of $Z$. Then as the coefficients of $D_{Y^{\prime}}$ are $\leq 1$, we have $D_{Y^{\prime}}\leq D^{\prime}$ over the generic point of $Z$. On the other hand, as the support of the fibre of $g^{\prime}$ over $z\in Z$ is contained in the toroidal boundary $D^{\prime}$, shrinking $Z$ around $z$ if necessary, we can assume that $D_{Y^{\prime}}\leq D^{\prime}$ over the whole $Z$.
@@
[…]
Pick an irreducible component of $S$ containing $V$ and let $T$ be its normalisation. Applying [19, Theorem 1.1], we can write $(K_{Y^{\prime}}+D^{\prime})|_{T}=K_{T}+D_{T}$, where $(T,D_{T})$ is lc. (Since $\dim T=\dim Y^{\prime}-1$, there is no nef part of the generalised pair in this adjunction; see [19, Corollary 1.4].) Take a general Cartier divisor $P\geq 0$ on $Y^{\prime}$ that contains $V$ and does not contain other lc centres of $(Y^{\prime},D^{\prime})$ lying outside $V$. For any small $\alpha>0$, $(Y^{\prime},D^{\prime}+\alpha P)$ is not lc near $V$, hence applying [19, Theorem 1.1] again, we deduce that $(T,D_{T})$ has an lc centre $V_{T}$ mapping onto $V$. Indeed, by [19, Theorem 1.1], we can write
$(K_{Y^{\prime}}+D^{\prime}+\alpha P)|_{T}=K_{T}+D_{T}+\alpha P_{T},$
where $P_{T}$ is the pullback of $P$ to $T$, and $(T,D_{T}+\alpha P_{T})$ is not lc for any $\alpha>0$. Then as $(T,D_{T})$ is lc, there exists an lc centre $Q_{T}$ of $(T,D_{T})$ contained in $\operatorname{Supp}P_{T}$. Taking $P$ generally, we see that the image of $Q_{T}$ in $Y^{\prime}$ is contained in $V$. Assume that no such lc centre $Q_{T}$ of $(T,D_{T})$ is mapped onto $V$. Then there is a proper closed subset $\Gamma\subset V$ such that the image of each such $Q_{T}$ in $Y^{\prime}$ is contained in $\Gamma$. Applying [19, Theorem 1.1] and the argument as above to the pair $(Y^{\prime}\setminus\Gamma,D^{\prime}|_{Y^{\prime}\setminus\Gamma})$, we see that there exists an lc centre of $(T,D_{T})$ whose generic point has image in $Y^{\prime}$ contained in $V\setminus\Gamma$. This is a contradiction by construction of $\Gamma\subset V$, hence there is an lc centre $V_{T}$ of $(T,D_{T})$ mapping onto $V$.
[…]
Step 6. Now we can conclude the proof as follows. Recall that $\operatorname{centre}_{Y^{\prime}}(F)$ is an lc centre of the pair $(Y^{\prime},D^{\prime})$ that is contained in $S$. By Step 5, $\operatorname{centre}_{Y^{\prime}}(F)$ is birationally bounded. Denote by $C$ a bounded nonsingular resolution of $\operatorname{centre}_{Y^{\prime}}(F)$. Hence the irrationality $\operatorname{irr}(C)$ is bounded from above by Lemma 2.17. Moreover, by Proposition 3.5, $F\dashrightarrow\operatorname{centre}_{Y^{\prime}}(F)$ has irreducible and rational general closed fibres, so does the rational map $F\dashrightarrow C$. ∎
@@
Let $F$ be an irreducible component of a closed fibre of $f$. Then there is a dominant rational map $F\dashrightarrow C$ whose general fibres are irreducible and rational, and $C$ is a bounded smooth projective variety depending only on $d,\epsilon$, hence with bounded degree of irrationality.
@@
Up to decreasing $\epsilon$ slightly, we can take an effective $\mathbb{Q}$-divisor $B$ so that $(X,B)$ is $\epsilon$-lc, and $K_{X}+B\sim_{\mathbb{Q}}0/Z$. By taking a $\mathbb{Q}$-factorial dlt model of $(X,B)$, we can assume that $X$ is $\mathbb{Q}$-factorial, and that the contraction $f\colon X\to Z$ is of Fano type; see [7, §2.13 (7)]. Then by [12], we can run MMP over $Z$ on any $\mathbb{R}$-divisor on $X$.
Let $z\in Z$ be the closed point such that $F$ is contained in the fibre $f^{-1}(z)$. By [12], we can run a (F)(-F)-MMP $\phi\colon X\dashrightarrow X^{\prime}$ over $Z$ that ends with a good minimal model of F-F. By negativity lemma, $F$ is not contracted by $\phi$. Let $F^{\prime}\coloneqq\phi_{*}F$. Then $F^{\prime}$ is supported on the whole fibre of $X^{\prime}\to Z$ over $z$. As $F$ is vertical/Z/Z, $\phi$ only modifies the fibre $f^{-1}(z)$, and leaves everything outside $f^{-1}(z)$ unchanged, so $K_{X^{\prime}}$ is big/Z/Z. Again by [12], we can run a $(-K_{X^{\prime}})$-MMP over $Z$, which induces a Fano fibration $g\colon X^{\prime\prime}\to Z$. Denote by $F^{\prime\prime}$ the pushdown of $F^{\prime}$ to $X^{\prime\prime}$. As $F^{\prime}$ is supported on the whole fibre of $X^{\prime}\to Z$ over $z$, we see that $F^{\prime\prime}$ is an irreducible divisor supported on the whole fibre of $g$ over $z$.
Let $B^{\prime\prime}$ be the pushdown of $B$ to $X^{\prime\prime}$. Then $(X^{\prime\prime},B^{\prime\prime})$ is also $\epsilon$-lc, and $K_{X^{\prime\prime}}+B^{\prime\prime}\sim_{\mathbb{Q}}0/Z$. By the canonical bundle formula [11, §2.11], we can write $K_{X^{\prime\prime}}+B^{\prime\prime}\sim_{\mathbb{Q}}g^{*}(K_{Z}+B_{Z}+M_{Z})$, where $(Z,B_{Z}+M_{Z})$ is a generalised pair. By [10, Theorem 1.1], there is a $\delta>0$ depending only on $d,\epsilon$ such that $(Z,B_{Z}+M_{Z})$ is generalised $\delta$-lc. Then the coefficient of $z$ in $B_{Z}$ is $\leq 1-\delta$. Thus, the lc threshold $\alpha$ of $g^{*}z$ with respect to $(X^{\prime\prime},B^{\prime\prime})$ is bounded from below: indeed, the coefficient of $z$ in $B_{Z}$ is $1-\alpha$, which implies that $\alpha\geq\delta$. Then there exists a $t>0$ depending only on $d,\epsilon$ such that $(X^{\prime\prime},B^{\prime\prime}+tF^{\prime\prime})$ is lc, hence $(X^{\prime\prime},tF^{\prime\prime})$ is lc. Thus, we can conclude by Theorem 1.3. ∎
Caucher Birkar
Yau Mathematical Sciences Center, Tsinghua University, Beijing, China
Email: birkar@mail.tsinghua.edu.cn
Santai Qu
Institute of Geometry and Physics, University of Science and Technology of China, Hefei, Anhui Province, China
Email: santaiqu@ustc.edu.cn